3. Consider the following table: 1995 S10.00 100 1985 2014 S20.00 200 Average wage S7 S7.00 CPI (base 1995) 80 Do these problems i. Assume 1995 is the base year, convert the nominal wages to constant dollar wages Assume 2014 is the base year, convert the nominal wages to constant dollar wages ii.
Answer #1

1-Base year-1995

a) For 1985, constant dollar wages= nominal wage/price index=7/(80/100)=8.75$

b) For 1995, constant dollar wages= nominal wage/price index=10/(100/100)=10$

c) For 2014, constant dollar wages= nominal wage/price index=20/(200/100)=10$

2- Base year- 2014

a) For 1985, constant dollar wages= nominal wage/ price index=7/(80/200)=17.5$

b) For 1995, constant dollar wages= nominal wage/price index=10/(100/200)=20$

c) For 2014, constant dollar wages= nominal wages/ price index=20/(200/200)= 20$